Hi. I was wondering if you could help me with this problem.
Given that f(3)=12 and f(5)=-2 find f(1/7)
I've tried the f(x)=mx+b formula
f(3)=m(3)+b=12
f(5)=m(5)+b=-2
5-3=2m and -2-12=-14
M=-14/-2
M=-7
And this is where I get stuck.
Thank you

Answer by jsmallt9(3759)

>Given that f(3)=12 and f(5)=-2 find f(1/7)
>I've tried the f(x)=mx+b formula
Good idea.
>f(3)=m(3)+b=12
>f(5)=m(5)+b=-2
The way you've written this is confusing. It would be better as:
12 = m(3) + b
-2 = m(5) + b
>M=-14/-2
>M=-7
Still good.

Now we just use this value for "m" to find "b". We can substitute into either equation:
12 = (-7)(3) + b
12 = -21 + b
33 = b

So the function is:
f(x) = -7x + 33

And finally we can use this to find f(1/7):
f(1/7) = -7(1/7) + 33
f(1/7) = -1 + 33
f(1/7) = 32